OpenStudy (anonymous):

If BC = 2x + 1, CD = 3x - 4 and BD = 22, find the value of x.

OpenStudy (anonymous):

BD=BC+CD

OpenStudy (anonymous):

CAN YOU SUBSTITUTE

OpenStudy (anonymous):

yes so it would be (2x-1) + (3x-4)

OpenStudy (anonymous):

=22

OpenStudy (anonymous):

\[22=2x+1+3x-4\]

OpenStudy (anonymous):

5x + 3 = 22 ?

OpenStudy (anonymous):

It would be 5x-3=22

OpenStudy (anonymous):

@Ivasquev462 used 2x-1 instead of 2x+1

OpenStudy (anonymous):

@lvasquev462

OpenStudy (anonymous):

ohh ok so

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

yes x is 5